OpenStudy (anonymous):

write 4km2/12k in simplest form

OpenStudy (anonymous):

simplify top and bottom by the common factor.

OpenStudy (anonymous):

what is the common factor?

OpenStudy (campbell_st):

I'll rewrite the expression and hopefully you'll be able to cancel common factors.. \[\frac{ 4 \times k \times m^2}{3 \times 4 \times k}\] anything that is common in the numerator and denominator can be cancelled... hope it helps

OpenStudy (unklerhaukus):

\[\frac{ \cancel{4 \times k \times}\quad m^2}{3\quad \cancel{\times 4 \times k}}\]